OpenStudy (anonymous):

The cost of a ticket to an amusement park was increased by . The new cost is . a. Write a one-step equation that could be solved to find the old price. Let p represent the old price. b. Solve the equation you wrote in Part a. Show all your work. c. Explain how you know that the solution you found in Part b is the correct answer.

OpenStudy (anonymous):

what's the new cost? I can't see

OpenStudy (anonymous):

oh sorry The cost of a ticket to an amusement park was increased by 4 dollars. The new cost is twenty-eight dollars. a. Write a one-step equation that could be solved to find the old price. Let p represent the old price. b. Solve the equation you wrote in Part a. Show all your work. c. Explain how you know that the solution you found in Part b is the correct answer.

OpenStudy (anonymous):

there @Owl_Eyes03 and @ArtemisWolf

OpenStudy (anonymous):

A) p + 4 = 28 B) 24 C) I know because I did the problem p + 4 - 4 = p = 28 - 4 = 24

OpenStudy (anonymous):

go with ArtemisWolf's :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

loll thanks

OpenStudy (anonymous):

No prob